OpenStudy (anonymous):

If y varies directly as x and y = 14 when x = 8, find x when y = 21. A. 24 B. 12 C. 36 D. 16

OpenStudy (anonymous):

14 : 8 21 : x cross multiply

OpenStudy (anonymous):

i got 1.14

OpenStudy (anonymous):

then you did not do it correctly

OpenStudy (anonymous):

14x=8 *21 x = 21 *8/14

OpenStudy (anonymous):

its b ... sry im acting dumb :(:(.. its like 3 in the morning and if i dont finish my work im going to fail :(

OpenStudy (anonymous):

im sry :(

OpenStudy (anonymous):

hey, no problem. just try to keep it straight in the calculator, otherwise you will be prone to making mistakes. brackets are important.

OpenStudy (anonymous):

ok i will.